hi, can you help me solve this twp step equation: 16n+2= -46
Answer by mananth(16946)

16n+2=-46
add =-2
16n+2-2=-46-2
-16n = -48
/-16
-16n/-16 = -48/-16
n= 3
...
m.ananth@hotmail.ca